Grade 11Physical Chemistry

A 5.2 molal aqueous solution of methyl alcohol, 3 CH OH, is supplied. What is the mole fraction of methyl alcohol in the solution ?

tanisha singh
12 Years agoGrade 11

1 Answer

Rijit Singh
10 Years ago
Molarity= moles of solute/mass of solvent=5.2
 
moles of solute= 5.2
mass of solvent=1kg=1000g
 
no of moles=given mass / molar mass
 
molar mass of water is 18.
 
then no. of moles of solvent or water is 1000/18=55.55
 
molar fraction= moles of solute/moles of solute+moles of solvent
 
molar fravtion=5.5/5.5+55.55
molar fraction=5.5/60.6